# EMMA-500: Enhancing Massively Multilingual Adaptation of Large Language Models

### v1

@@ D.4 Math
Llama 2 7B 0.0669 0.0280 0.0105 0.0800 0.0172 0.1760 0.0241 0.1120 0.0200 0.1200 0.0206 0.0240 0.0097 0.0800 0.0172 0.0280 0.0105 0.0120 0.0069 0.0080 0.0056 0.0680 0.0160
Llama 2 7B Chat 0.1022 0.0280 0.0105 0.1680 0.0237 0.2280 0.0266 0.1960 0.0252 0.1920 0.0250 0.0240 0.0097 0.1440 0.0222 0.0080 0.0056 0.0080 0.0056 0.0280 0.0105 0.1000 0.0190
CodeLlama 2 7B 0.0593 0.0160 0.0080 0.0880 0.0180 0.1280 0.0212 0.0880 0.0180 0.1040 0.0193 0.0560 0.0146 0.0600 0.0151 0.0200 0.0089 0.0120 0.0069 0.0520 0.0141 0.0280 0.0105
LLaMAX Llama 2 7B 0.0335 0.0280 0.0105 0.0360 0.0118 0.0600 0.0151 0.0200 0.0089 0.0720 0.0164 0.0320 0.0112 0.0240 0.0097 0.0160 0.0080 0.0080 0.0056 0.0160 0.0080 0.0560 0.0146
LLaMAX Llama 2 7B Alpaca 0.0505 0.0400 0.0124 0.0360 0.0118 0.1080 0.0197 0.0600 0.0151 0.0640 0.0155 0.0320 0.0112 0.0440 0.0130 0.0480 0.0135 0.0160 0.0080 0.0320 0.0112 0.0760 0.0168
MaLA-500 Llama 2 10B v1 0.0091 0.0000 0.0000 0.0000 0.0000 0.0120 0.0069 0.0200 0.0089 0.0240 0.0097 0.0120 0.0069 0.0200 0.0089 0.0040 0.0040 0.0000 0.0000 0.0040 0.0040 0.0040 0.0040
MaLA-500 Llama 2 10B v2 0.0091 0.0000 0.0000 0.0000 0.0000 0.0120 0.0069 0.0200 0.0089 0.0240 0.0097 0.0120 0.0069 0.0200 0.0089 0.0040 0.0040 0.0000 0.0000 0.0040 0.0040 0.0040 0.0040
YaYi Llama 2 7B 0.0709 0.0320 0.0112 0.0840 0.0176 0.1560 0.0230 0.1600 0.0232 0.1040 0.0193 0.0520 0.0141 0.0560 0.0146 0.0080 0.0056 0.0120 0.0069 0.0040 0.0040 0.1120 0.0200
TowerBase Llama 2 7B 0.0615 0.0240 0.0097 0.0840 0.0176 0.1160 0.0203 0.0920 0.0183 0.0880 0.0180 0.0480 0.0135 0.1000 0.0190 0.0120 0.0069 0.0080 0.0056 0.0160 0.0080 0.0880 0.0180
TowerInstruct Llama 2 7B 0.0724 0.0160 0.0080 0.1000 0.0190 0.1520 0.0228 0.1560 0.0230 0.1280 0.0212 0.0160 0.0080 0.1000 0.0190 0.0160 0.0080 0.0200 0.0089 0.0200 0.0089 0.0720 0.0164
Occiglot Mistral 7B v0.1 0.1331 0.0320 0.0112 0.2120 0.0259 0.3000 0.0290 0.2720 0.0282 0.2160 0.0261 0.0640 0.0155 0.1520 0.0228 0.0240 0.0097 0.0160 0.0080 0.0800 0.0172 0.0960 0.0187
Occiglot Mistral 7B v0.1 Instruct 0.2276 0.0480 0.0135 0.3400 0.0300 0.4640 0.0316 0.4000 0.0310 0.3160 0.0295 0.1840 0.0246 0.2360 0.0269 0.0640 0.0155 0.0280 0.0105 0.1120 0.0200 0.3120 0.0294
BLOOM 7B 0.0287 0.0240 0.0097 0.0160 0.0080 0.0400 0.0124 0.0360 0.0118 0.0120 0.0069 0.0200 0.0089 0.0360 0.0118 0.0400 0.0124 0.0360 0.0118 0.0200 0.0089 0.0360 0.0118
BLOOMZ 7B 0.0255 0.0320 0.0112 0.0160 0.0080 0.0360 0.0118 0.0240 0.0097 0.0320 0.0112 0.0280 0.0105 0.0360 0.0118 0.0280 0.0105 0.0120 0.0069 0.0120 0.0069 0.0240 0.0097
mGPT 0.0135 0.0000 0.0000 0.0200 0.0089 0.0320 0.0112 0.0080 0.0056 0.0240 0.0097 0.0160 0.0080 0.0240 0.0097 0.0160 0.0080 0.0000 0.0000 0.0000 0.0000 0.0080 0.0056
mGPT 13B 0.0131 0.0000 0.0000 0.0160 0.0080 0.0160 0.0080 0.0080 0.0056 0.0360 0.0118 0.0160 0.0080 0.0200 0.0089 0.0120 0.0069 0.0000 0.0000 0.0080 0.0056 0.0120 0.0069
XGLM 7.5B 0.0102 0.0000 0.0000 0.0080 0.0056 0.0000 0.0000 0.0120 0.0069 0.0000 0.0000 0.0040 0.0040 0.0200 0.0089 0.0200 0.0089 0.0200 0.0089 0.0000 0.0000 0.0280 0.0105
YaYi 7B 0.0276 0.0240 0.0097 0.0200 0.0089 0.0600 0.0151 0.0240 0.0097 0.0560 0.0146 0.0160 0.0080 0.0120 0.0069 0.0120 0.0069 0.0240 0.0097 0.0080 0.0056 0.0480 0.0135
Llama 3 8B 0.2745 0.1760 0.0241 0.3960 0.0310 0.5080 0.0317 0.4720 0.0316 0.3640 0.0305 0.0360 0.0118 0.3760 0.0307 0.2400 0.0271 0.0560 0.0146 0.3680 0.0306 0.0280 0.0105
Llama 3.1 8B 0.2836 0.2000 0.0253 0.4120 0.0312 0.5520 0.0315 0.4840 0.0317 0.3840 0.0308 0.0360 0.0118 0.4000 0.0310 0.2080 0.0257 0.0440 0.0130 0.3760 0.0307 0.0240 0.0097
Gemma 7B 0.3822 0.3440 0.0301 0.4480 0.0315 0.5880 0.0312 0.4800 0.0317 0.3960 0.0310 0.1680 0.0237 0.4120 0.0312 0.3760 0.0307 0.2720 0.0282 0.4280 0.0314 0.2920 0.0288
Gemma 2 9B 0.3295 0.2960 0.0289 0.4040 0.0311 0.5640 0.0314 0.5080 0.0317 0.3600 0.0304 0.0120 0.0069 0.3520 0.0303 0.3760 0.0307 0.3080 0.0293 0.4000 0.0310 0.0440 0.0130
Qwen 2 7B 0.4895 0.4440 0.0315 0.6560 0.0301 0.8080 0.0250 0.7600 0.0271 0.6960 0.0292 0.0160 0.0080 0.6720 0.0298 0.1600 0.0232 0.1480 0.0225 0.5600 0.0315 0.4640 0.0316
Qwen 1.5 7B 0.3156 0.1240 0.0209 0.4400 0.0315 0.5520 0.0315 0.4840 0.0317 0.4520 0.0315 0.1680 0.0237 0.4320 0.0314 0.0720 0.0164 0.0400 0.0124 0.2840 0.0286 0.4240 0.0313
EMMA-500 Llama 2 7B 0.1702 0.0880 0.0180 0.2320 0.0268 0.3400 0.0300 0.2800 0.0285 0.2560 0.0277 0.0920 0.0183 0.2280 0.0266 0.1680 0.0237 0.0240 0.0097 0.1000 0.0190 0.0640 0.0155
@@
Table 29: 3-shot results (ACC) on MGSM by CoT prompting and flexible matching.

### v2

@@ Appendix E Detailed Results
Llama 2 7B 6.69 2.80 1.05 8.00 1.72 17.60 2.41 11.20 2.00 12.00 2.06 2.40 0.97 8.00 1.72 2.80 1.05 1.20 0.69 0.80 0.56 6.80 1.60
Llama 2 7B Chat 10.22 2.80 1.05 16.80 2.37 22.80 2.66 19.60 2.52 19.20 2.50 2.40 0.97 14.40 2.22 0.80 0.56 0.80 0.56 2.80 1.05 10.00 1.90
CodeLlama 2 7B 5.93 1.60 0.80 8.80 1.80 12.80 2.12 8.80 1.80 10.40 1.93 5.60 1.46 6.00 1.51 2.00 0.89 1.20 0.69 5.20 1.41 2.80 1.05
LLaMAX Llama 2 7B 3.35 2.80 1.05 3.60 1.18 6.00 1.51 2.00 0.89 7.20 1.64 3.20 1.12 2.40 0.97 1.60 0.80 0.80 0.56 1.60 0.80 5.60 1.46
LLaMAX Llama 2 7B Alpaca 5.05 4.00 1.24 3.60 1.18 10.80 1.97 6.00 1.51 6.40 1.55 3.20 1.12 4.40 1.30 4.80 1.35 1.60 0.80 3.20 1.12 7.60 1.68
MaLA-500 Llama 2 10B v1 0.91 0.00 0.00 0.00 0.00 1.20 0.69 2.00 0.89 2.40 0.97 1.20 0.69 2.00 0.89 0.40 0.40 0.00 0.00 0.40 0.40 0.40 0.40
MaLA-500 Llama 2 10B v2 0.91 0.00 0.00 0.00 0.00 1.20 0.69 2.00 0.89 2.40 0.97 1.20 0.69 2.00 0.89 0.40 0.40 0.00 0.00 0.40 0.40 0.40 0.40
YaYi Llama 2 7B 7.09 3.20 1.12 8.40 1.76 15.60 2.30 16.00 2.32 10.40 1.93 5.20 1.41 5.60 1.46 0.80 0.56 1.20 0.69 0.40 0.40 11.20 2.00
TowerBase Llama 2 7B 6.15 2.40 0.97 8.40 1.76 11.60 2.03 9.20 1.83 8.80 1.80 4.80 1.35 10.00 1.90 1.20 0.69 0.80 0.56 1.60 0.80 8.80 1.80
TowerInstruct Llama 2 7B 7.24 1.60 0.80 10.00 1.90 15.20 2.28 15.60 2.30 12.80 2.12 1.60 0.80 10.00 1.90 1.60 0.80 2.00 0.89 2.00 0.89 7.20 1.64
Occiglot Mistral 7B v0.1 13.31 3.20 1.12 21.20 2.59 30.00 2.90 27.20 2.82 21.60 2.61 6.40 1.55 15.20 2.28 2.40 0.97 1.60 0.80 8.00 1.72 9.60 1.87
Occiglot Mistral 7B v0.1 Instruct 22.76 4.80 1.35 34.00 3.00 46.40 3.16 40.00 3.10 31.60 2.95 18.40 2.46 23.60 2.69 6.40 1.55 2.80 1.05 11.20 2.00 31.20 2.94
BLOOM 7B 2.87 2.40 0.97 1.60 0.80 4.00 1.24 3.60 1.18 1.20 0.69 2.00 0.89 3.60 1.18 4.00 1.24 3.60 1.18 2.00 0.89 3.60 1.18
BLOOMZ 7B 2.55 3.20 1.12 1.60 0.80 3.60 1.18 2.40 0.97 3.20 1.12 2.80 1.05 3.60 1.18 2.80 1.05 1.20 0.69 1.20 0.69 2.40 0.97
mGPT 1.35 0.00 0.00 2.00 0.89 3.20 1.12 0.80 0.56 2.40 0.97 1.60 0.80 2.40 0.97 1.60 0.80 0.00 0.00 0.00 0.00 0.80 0.56
mGPT 13B 1.31 0.00 0.00 1.60 0.80 1.60 0.80 0.80 0.56 3.60 1.18 1.60 0.80 2.00 0.89 1.20 0.69 0.00 0.00 0.80 0.56 1.20 0.69
YaYi 7B 2.76 2.40 0.97 2.00 0.89 6.00 1.51 2.40 0.97 5.60 1.46 1.60 0.80 1.20 0.69 1.20 0.69 2.40 0.97 0.80 0.56 4.80 1.35
Aya 23 8B 22.29 3.20 1.12 37.20 3.06 50.00 3.17 41.20 3.12 39.20 3.09 27.60 2.83 34.00 3.00 2.80 1.05 0.80 0.56 4.40 1.30 4.80 1.35
Aya expanse 8B 43.02 22.40 2.64 69.20 2.93 78.40 2.61 72.00 2.85 63.60 3.05 50.80 3.17 69.60 2.92 12.80 2.12 5.20 1.41 21.60 2.61 7.60 1.68
Llama 3 8B 27.45 17.60 2.41 39.60 3.10 50.80 3.17 47.20 3.16 36.40 3.05 3.60 1.18 37.60 3.07 24.00 2.71 5.60 1.46 36.80 3.06 2.80 1.05
Llama 3.1 8B 28.36 20.00 2.53 41.20 3.12 55.20 3.15 48.40 3.17 38.40 3.08 3.60 1.18 40.00 3.10 20.80 2.57 4.40 1.30 37.60 3.07 2.40 0.97
Gemma 7B 38.22 34.40 3.01 44.80 3.15 58.80 3.12 48.00 3.17 39.60 3.10 16.80 2.37 41.20 3.12 37.60 3.07 27.20 2.82 42.80 3.14 29.20 2.88
Gemma 2 9B 32.95 29.60 2.89 40.40 3.11 56.40 3.14 50.80 3.17 36.00 3.04 1.20 0.69 35.20 3.03 37.60 3.07 30.80 2.93 40.00 3.10 4.40 1.30
Qwen 2 7B 48.95 44.40 3.15 65.60 3.01 80.80 2.50 76.00 2.71 69.60 2.92 1.60 0.80 67.20 2.98 16.00 2.32 14.80 2.25 56.00 3.15 46.40 3.16
Qwen 1.5 7B 31.56 12.40 2.09 44.00 3.15 55.20 3.15 48.40 3.17 45.20 3.15 16.80 2.37 43.20 3.14 7.20 1.64 4.00 1.24 28.40 2.86 42.40 3.13
EMMA-500 Llama 2 7B 17.02 8.80 1.80 23.20 2.68 34.00 3.00 28.00 2.85 25.60 2.77 9.20 1.83 22.80 2.66 16.80 2.37 2.40 0.97 10.00 1.90 6.40 1.55
@@
Table 26: 3-shot results (ACC %) on MGSM in all languages by CoT prompting and flexible matching.